7. Consider the IR, H-NMR and C-NMR spectra of the compound with MFC,1,0, Identify the structure of the unknown compound MP8. Consider the IR, H-NMR and 13C-NMR spectra of the compound with MF: C,H, Bro, Identify the structure of the unknown compoCopyright © 1994 HAVENUMBERS 11.6 PPM 11.6 PPM Proton NMA 5 .5*3.6 2.2.0 .5 PPX 2.0 А0 T T 20 PPM 09 TTTTTTTTTTTTTTTTTTTTTTTT
Answer #1

7) C7H140: Degree of unsaturation ((2C+2)+N-H-X)/2 = (16-14)/2= 1 The IR peak at 1712 indicates the presence of carbonyl and